# Accelerating Diffusion Transformers with Dual Feature Caching

### v1

@@ 3.2 Dual Feature Caching
$\mathcal{G}(x)_{t}:=g_{l+1}\circ...\circ g_{L}^{t}(\mathcal{C}[l]).$ (3)
To achieve the maximal acceleration with aggressive caching, we set $l=L-1$ in most of our experiments, which indicates almost skipping all the layers in the caching step.
@@
As shown in Figure 1(a), conservative caching does not cache and reuse all the features. Instead, it caches and reuses only the features without the residual connection, which allows it to allocate the computation to some layers in the caching steps to correct the cached features. Recently, ToCa [35] introduced a token-wise approach that selectively reuses features based on the importance of each token and its sensitivity to caching, which enables us to perform computation only in the important tokens. Following ToCa, we first calculate an importance score $\mathcal{S}$ for each token $x_{i}$ and then use this score to divide the index set of tokens $\mathcal{I}$ into two complementary sets: $\mathcal{I}_{\text{Cache}}$, for tokens to be cached, and $\mathcal{I}_{\text{Compute}}$, for tokens to be computed. Subsequently, the computation for the $i_{th}$ token $x_{i}$ on the $l_{th}$ layer can be expressed as:
$\gamma_{i}f^{l}(x_{i})+(1-\gamma_{i})\mathcal{C}[l](x_{i}),$ (4)
where $\gamma_{i}=0$ for $i\in\mathcal{I}_{\text{Cache}}$ and $\gamma_{i}=1$ for $i\in\mathcal{I}_{\text{Compute}}$. Here, $\mathcal{C}[l](x_{i})$ denotes retrieving the cached value of $x_{i}$ from the conservative cache $\mathcal{C}$ at layer $l$, which incurs no computation cost, enabling partial omission of calculations. After this selective computation in the token dimension, the computed features are updated to $\mathcal{C}[l]$ for future reusing.
Value-based Token-wise Feature Caching:
@@
As the previous token-wise conservative caching method, ToCa offers two attention map-based selection strategies that have shown good effectiveness. However, the attention map-based selection strategy is incompatible with another commonly used attention-optimization method FlashAttention[9] and memory-efficient attention, which can not provide the attention scores but are capable of reducing the memory costs of attention from $O(N^{2})$ to $O(N)$ and significant accelerating the attention layers. This conflict between attention scores-based token selection with this optimized attention computation makes ToCa less practical in application and raises the requirements for attention score-free token selection methods.
Figure 4 shows the relation between the norm of the value matrix and the attention scores of each token, which demonstrates that tokens with high attention scores in the self-attention map tend to have relatively smaller norms in their value matrix. This suggests that we can directly select tokens with smaller norms in their value matrix as substitutes for those with high attention map scores. Concretely, the value matrix of each token is produced by a linear layer in attention layers as follows: $v_{i}=\text{MLP}_{v}(x_{i})$, where $v_{i}$ is value vector for the $i^{th}$ token. Therefore, V-Caching gives the selection rule for $\mathcal{I}_{\text{Cache}}$ as follows:
$\mathcal{I}_{\text{Cache}}=\mathop{\arg\max}_{\{i_{1},i_{2},\dots,i_{n}\}\subseteq\{1,2,\dots,N\}}\{\|v_{i_{1}}\|_{2},\|v_{i_{2}}\|_{2},\cdots,\|v_{i_{n}}\|_{2}\},$ (5)
where the $n$ is the total number of selected tokens, defined by $n=\lfloor R\%\times N\rfloor$.
3.3 Overall Framework
@@
Built on the deeper understanding of aggressive and conservative caching, DuCa introduces a caching scheme that alternates conservative and aggressive caching in different steps. V-Caching serves as the conservative caching strategy, as shown in Figure 3. This design leverages the realignment capability of token-wise conservative caching to correct the misaligned timestep information occurring in aggressive caching steps. Therefore, it enables the high local compression rate of aggressive caching to be applied to more steps while further reducing redundant computations in conservative caching steps. In practice, we arrange token-wise conservative caching steps on odd-numbered steps following a fresh step, and aggressive caching steps on even-numbered steps.
Figure 4: The relation between the self-attention score and the norm of the value matrix for each token. Tokens with the largest attention score tend to have a lower norm in the value matrix.
4 Experiments
4.1 Experiment Settings
4.1.1 Model Configurations
Experiments are conducted on three widely known DiT-based models on the three generation tasks with the correlating default samplers, including PixArt-$\alpha$ [7] with 20 DPM Solver++ [23] steps for text-to-image generation, OpenSora [34] with 30 reflected flow steps [21] for text-to-video generation, and DiT-XL/2 [26] with 50 DDIM [31] steps for class-conditional image generation. All the experiments are conducted with NVIDIA H800 80GB GPUs. See More details in our appendix.
4.1.2 Evaluation and Metrics
Text-to-image generation We randomly select 30,000 captions from COCO-2017 datasets [20] to generate the 30,000 $256\times 256$ images. FID-30k is computed using the 30k generated images and the real images from COCO-2017 to evaluate the quality of text-to-image generation. Besides, the CLIP score [14] is computed between the 30k prompt-image pairs to evaluate the text-image alignment. All the CLIP scores are tested using the Vit-large-14 model.
Text-to-video generation VBench [17] evaluation framework is applied to evaluate the performance of the text-to-video model. With 5 videos per prompt and 950 prompts from the framework, 4,750 videos are generated to evaluate the 16 aspects proposed by VBench. The generated videos are 480p, 9:16, and 2s videos.
Class-conditional image generation 50,000 $256\times 256$ images uniformly sampled from the 1,000 classes in ImageNet[10] are generated to evaluate the FID-50k [15]. Besides, sFID, Precision, and Recall are introduced as supplementary metrics.
@@ 4.1.2 Evaluation and Metrics
Table 1: Quantitative comparison of text-to-image generation on MS-COC02017 with PixArt-$\alpha$ and 20 DPM++ sampling steps.
@@
PixArt-$\alpha$ [7] 0.242 11.88 1.00$\times$ 28.12 16.29
$50\%$ steps 0.121 5.94 2.00$\times$ 37.57 15.83
FORA [28] 0.178 5.66 1.98$\times$ 29.67 16.40
ToCa[35] 0.167 6.05 1.96$\times$ 28.38 16.43
DuCa(C-A score) 0.157 5.99 1.98$\times$ 27.98 16.44
DuCa(V-Caching) 0.088 5.99 1.98$\times$ 28.05 16.42
Table 2: Quantitative comparison in text-to-video generation on VBench. *Results are from PAB [33]. PAB1-3 indicate PAB with different hyper-parameters.
[…]
OpenSora [34] 54.83 3283.20 1.00$\times$ 79.41
@@
PAB1∗ [33] 41.05 2657.70 1.24$\times$ 78.51
PAB2∗ [33] 39.96 2615.15 1.26$\times$ 77.64
PAB3∗ [33] 38.26 2558.25 1.28$\times$ 76.95
$50\%$ steps 28.85 1641.60 2.00$\times$ 76.78
FORA [28] 33.19 1751.32 1.87$\times$ 76.91
ToCa[35] 29.03 1394.03 2.36$\times$ 78.59
DuCa(C-A score) 26.77 1315.62 2.50$\times$ 78.63
DuCa(V-Caching) 19.16 1315.62 2.50$\times$ 78.54
Table 3: Quantitative comparison on class-to-image generation on ImageNet with DiT-XL/2.
@@
DiT-XL/2-G (cfg $=1.50$) 2.012 118.68 1.00$\times$ 4.98 2.31 0.82 0.58
DDIM-50 steps 0.417 23.74 1.00$\times$ 4.40 2.43 0.80 0.59
DDIM-25 steps 0.209 11.87 2.00$\times$ 4.74 3.18 0.79 0.58
DDIM-20 steps 0.167 9.49 2.50$\times$ 5.15 3.81 0.78 0.58
DDIM-17 steps 0.142 8.07 2.94$\times$ 5.76 4.58 0.77 0.56
FORA [28] 0.197 4.29 2.77$\times$ 6.43 3.88 0.79 0.54
ToCa(a) 0.195 8.73 2.72$\times$ 5.14 3.64 0.79 0.55
ToCa(b) 0.214 10.23 2.32$\times$ 4.70 3.05 0.79 0.57
DuCa(a)(S-A score) 0.184 8.76 2.71$\times$ 4.66 3.19 0.79 0.57
DuCa(b)(S-A score) 0.200 9.58 2.48$\times$ 4.66 3.05 0.80 0.57
DuCa(a)(V-Caching) 0.177 8.76 2.71$\times$ 4.68 3.19 0.80 0.57
DuCa(b)(V-Caching) 0.198 9.58 2.48$\times$ 4.69 3.06 0.80 0.57
4.2 Quantitative and Qualitative Results
Figure 5: Visualization results for different acceleration methods on PixArt-$\alpha$.
@@ 4.2 Quantitative and Qualitative Results
4.2.1 Results on text-to-image generation
As shown in Table 1, we compared the DuCa under two different token selection methods: Cross-Attention score from ToCa and the proposed V-norm score with three training-free acceleration methods: ToCa[35], the 10 DPM++ sampling steps and FORA [28], a layer-wise conservative caching method.
In terms of generation quality, the quantitative results demonstrate that under an acceleration ratio close to 2.0, the DuCa methods achieve the lowest FID-30k score on the MS-COCO2017 generation task, with almost no loss in generation quality compared to the original non-accelerated approach. For instance, the DuCa method using the Cross-Attention score reduces the FID-30k by 0.40 compared to the ToCa method with the same acceleration ratio, while the DuCa method using the V-norm score also decreases the FID-30k by 0.33, demonstrating that our alternating acceleration strategy effectively enhances generation quality.
In terms of evaluating the text-image alignment metric, the performance of the DuCa method is very close to that of ToCa and generally superior to the approach of directly setting $50\%$ generation steps. For example, the DuCa method with Cross-Attention score and V-norm score achieves CLIP scores on MS-COCO2017 that are 0.61 and 0.59 higher, respectively, compared to the $50\%$ generation steps method with the same acceleration ratio. These scores are even slightly better than those of the original non-accelerated PixArt-$\alpha$ model, with improvements of 0.15 and 0.13, respectively, demonstrating better controllability.
In addition, thanks to the acceleration brought by the compatibility with FlashAttention [9], the latency of the DuCa method with V-norm score is reduced by 0.069 compared to the DuCa method with Attention score, representing an approximate $43.9\%$ reduction, which is a significant improvement in acceleration.
4.2.2 Results on text-to-video generation
Figure 7: VBench metrics and acceleration ratio of proposed DuCa and other methods
The generation results of $\Delta$-DiT, T-GATE[32], PAB[33], 15 rflow sampling steps, ToCa[35], FORA [28] and proposed DuCa on the text-to-video model OpenSora[34] are shown in Table 2.
As shown in Table 2, DuCa achieves a further acceleration from 2.36 $\times$ (as with ToCa) to 2.50 $\times$ with virtually no loss in generation quality. Additionally, despite having the lowest computational cost among all the compared methods, DuCa still achieves the highest VBench score. This demonstrates that the proposed strategy in DuCa, which alternates between aggressive and conservative caching steps, effectively eliminates the redundant components present in a fully conservative caching approach.
Additionally, comparing the generation results of the Cross-Attention and V-Caching token selection strategies, we only observe a slight decrease of 0.09 in the VBench score with the V-Caching strategy. More significantly, V-Caching reduces generation time by 7.61 seconds, approximately $28.4\%$. This indicates that DuCa effectively leverages V-Caching to make the acceleration method compatible with FlashAttention[9], achieving substantial reductions in generation time. In practice, using V-Caching will be more advantageous for rapid inference.
@@ 4.2.2 Results on text-to-video generation
In Figure 7, we further present the performance of DuCa and the aforementioned methods across various dimensions of the VBench metric. DuCa shows results consistent with the non-accelerated original OpenSora model across multiple metrics, including but not limited to Overall Consistency, Temporal Style, Scene, and Object Class. Additionally, the Cross-Attention and V-Caching strategies introduce only a slight difference in Spatial Relationship, further underscoring the exceptional performance of DuCa and the proposed V-Caching in practical scenarios.
4.2.3 Results on class-conditional image generation
@@
In Table 3, we compare ToCa[35], FORA [28], the direct reduction of corresponding DDIM steps, and the proposed DuCa under Self-Attention score and V-Caching token selection strategies. The experimental results indicate that with comparable FID scores, representing generation quality, DuCa further enhances acceleration. For instance, DuCa(b) with Self-Attention score achieves a similar FID-50k as ToCa(b), but the acceleration ratio increases from 2.36 to 2.48. Additionally, under a higher acceleration scenario with an acceleration ratio around 2.7, DuCa achieves an FID-50k of 3.19, further reducing the FID-50k by 0.45 compared to ToCa’s 3.64. This demonstrates that DuCa effectively reallocates computation to more critical areas.
4.3 Ablation Studies
Table 4: Ablation Study for different token selection methods for DuCa on PixArt-$\alpha$.
Methods MS-COCO2017 Latency(s) $\downarrow$ FLOPs(T) $\downarrow$
FID-30k $\downarrow$ CLIP $\uparrow$
Random 28.34 16.42 0.089 5.99
Cross-Attention 27.98 16.44 0.157 5.99
V-Caching 28.05 16.42 0.088 5.99
Table 5: Comparison on different caching methods on PixArt-$\alpha$.
Methods MS-COCO2017 FLOPs(T) $\downarrow$
FID-30k $\downarrow$ CLIP $\uparrow$
Conservative-only 27.92 16.39 7.84
Aggressive-only 30.51 16.39 3.59
DuCa 28.05 16.42 5.99
As shown in Table 4, we compare Random, Cross-Attention based, and V-Caching token selection methods on PixArt-$\alpha$ in terms of latency, FLOPs, FID-30k, and CLIP score. It can be observed that the cross-Attention-based selection method achieves the best performance in generation quality and text-image alignment, but due to its incompatibility with FlashAttention[9], its latency is as high as 0.157 seconds—nearly twice that of V-Caching and Random selection. In contrast, V-Caching achieves nearly the same, only slightly lower, FID-30k and CLIP scores, along with faster generation speed. Despite the minimal impact of the three token selection methods on computational load, the FlashAttention[9]-compatible V-Caching method shows significantly higher acceleration due to computational optimizations. Thus, V-Caching stands out as a superior generation scheme that effectively balances quality and efficiency.

### v2

@@ IV-B Dual Feature Caching
$\mathcal{G}(x)_{t}:=g_{l+1}\circ...\circ g_{L}^{t}(\mathcal{C}[l]).\vskip-8.53581pt$ (3)
To achieve the maximal acceleration with aggressive caching, we set l=L1l=L-1 in most of our experiments, which indicates almost skipping all the layers in the caching step.
@@
As shown in Figure 1(a), conservative caching does not cache and reuse all the features. Instead, it caches and reuses only the features without the residual connection, which allows it to allocate the computation to some layers in the caching steps to correct the cached features. Recently, ToCa [zou2024accelerating] introduced a token-wise approach that selectively reuses features based on the importance of each token and its sensitivity to caching, which enables us to perform computation only in the important tokens. Following ToCa, we first calculate an importance score $\mathcal{S}$ for each token $x_{i}$ and then use this score to divide the index set of tokens $\mathcal{I}$ into two complementary sets: $\mathcal{I}_{\text{Cache}}$, for tokens to be cached, and $\mathcal{I}_{\text{Compute}}$, for tokens to be computed. Subsequently, the computation for the $i_{th}$ token $x_{i}$ on the $l_{th}$ layer can be expressed as: $\gamma_{i}f^{l}(x_{i})+(1-\gamma_{i})\mathcal{C}[l](x_{i}),$ where $\gamma_{i}=0$ for $i\in\mathcal{I}_{\text{Cache}}$ and $\gamma_{i}=1$ for $i\in\mathcal{I}_{\text{Compute}}$. Here, $\mathcal{C}[l](x_{i})$ denotes retrieving the cached value of $x_{i}$ from the conservative cache $\mathcal{C}$ at layer $l$, which incurs no computation cost, enabling partial omission of calculations. After this selective computation in the token dimension, the computed features are updated to $\mathcal{C}[l]$ for future reusing.
Key Issue of Attention-Based Selection Methods
@@
As the previous token-wise conservative caching method, ToCa offers two attention map-based selection strategies that have shown good effectiveness. However, the attention map-based selection strategy is incompatible with another commonly used attention-optimization method FlashAttention[Dao2022FlashAttentionFA] and memory-efficient attention, which can not provide the attention scores but are capable of reducing the memory costs of attention from $O(N^{2})$ to $O(N)$ and significant accelerating the attention layers. This conflict between attention scores-based token selection with this optimized attention computation makes ToCa less practical in application and raises the requirements for attention score-free token selection methods.
TABLE I: Token selection methods for FLUX on Image Reward. The experiments were conducted on the FLUX.1-dev model with a full activation interval of 4, selecting $10\%$ of tokens at each cache step. A higher Image Reward means better performance.
Score Efficient Attention Max Min
Attention 0.9798 0.9707
K-norm 0.9783 0.9726
V-norm 0.9747 0.9798
Similarity 0.9715 0.9811
Random 0.9806
Original Model 0.9898
Diversity matters more in token selection
As shown in Table I, we compare attention scores with other classic token selection methods used in large language and multimodal models, such as K-norm and V-norm, which select tokens based on maximizing and minimizing the corresponding values, respectively. We then compare these results with the simplest and most naive random selection. Surprisingly, while the method of selecting tokens based on the highest attention score in ToCa is the best-performing approach among all standard-based selection methods, it still performs slightly worse than random selection.
The results suggest that rather than the traditional assumption that tokens can be distinguished as important or unimportant, selecting a diverse and broad range of tokens is more beneficial for performance. Based on this hypothesis, we designed an experiment that scores tokens based on similarity. For each cache step in Table I, $10\%$ of all tokens were selected. First, $1\%$ of all tokens were randomly chosen as the base tokens. Then, for the remaining $9\%$, tokens were selected based on two criteria: one method selects tokens with the highest similarity to the base tokens, aiming to choose tokens of the same type, while the other method selects those with the lowest similarity, encouraging greater diversity in the selected tokens. The method based on maximum similarity achieved the worst performance among all selection methods, while the method based on minimum similarity achieved the best performance, further confirming the importance of diversity in token selection. Considering that the difference between the minimum similarity-based method and the random selection method is small and it does not require additional computation, the random token selection method is recommended.
IV-C Overall Framework
@@
Based on the understanding of aggressive and conservative caching, DuCa introduces a caching scheme that alternates conservative and aggressive caching in different steps. Random selection serves as the selection strategy for conservative caching, as shown in Figure 3. This design leverages the realignment capability of token-wise conservative caching to correct the misaligned timestep information occurring in aggressive caching steps. Therefore, it enables the high local compression rate of aggressive caching to be applied to more steps while further reducing redundant computations in conservative caching steps. We arrange token-wise conservative caching steps on odd-numbered steps following a fresh step and aggressive caching steps on even-numbered steps.
@@
TABLE II: Quantitative comparison in text-to-image generation for FLUX on Image Reward.
[…]
FLUX.1[flux2024] Attention [Dao2022FlashAttentionFA] Latency(s) $\downarrow$ Speed $\uparrow$ FLOPs(T) $\downarrow$ Speed $\uparrow$ DrawBench Overall
@@
[…]
V Experiments
V-A Experiment Settings
V-A1 Model Configurations
Experiments are conducted on four widely known DiT-based models on the three generation tasks with the correlating default samplers, including the FLUX.1-dev model [flux2024] with 50 rectified flow sampling steps, HiDream model with 50 rectified flow sampling steps, OpenSora [opensora] with 30 rectified flow sampling steps [refitiedflow] for text-to-video generation, and DiT-XL/2 [peebles2023dit] with 50 DDIM [songDDIM] steps for class-conditional image generation. All the experiments are conducted with NVIDIA H800 80GB GPUs. Additionally, we have supplemented our experiments with the smaller text-to-image generation model PixArt-$\alpha$ [chen2023pixartalpha] with 20 DPM Solver++ [lu2022dpm++] steps and included the detailed design of all experiments. More details are available in the appendix.
V-A2 Evaluation and Metrics
Text-to-image generation For the FLUX.1-dev and HiDream model with a high resolution of $1000\times 1000$, we utilized 200 high-quality prompts from DrawBench [Saharia2022PhotorealisticTD], which focus on generating various types of images. We then generated images using these prompts and employed Image Reward [xuImageRewardLearningEvaluating2023] to assess the quality of the generated images while using the GenEval [Ghosh2023GenEvalAO] benchmarks to evaluate the alignment between the images and the text.
Text-to-video generation VBench [huang2024vbench] evaluation framework is applied to evaluate the performance of the text-to-video model. With 5 videos per prompt and 950 prompts from the framework, 4,750 480p, 9:16, and 2s videos are generated to evaluate the 16 aspects proposed by VBench.
TABLE III: Quantitative results for HiDream-l1.
Method Acceleration Image Reward $\uparrow$
HiDream-l1 Latency(s) $\downarrow$ Speed $\uparrow$ FLOPs(T) $\downarrow$ Speed $\uparrow$ DrawBench
[full]: 50 steps 48.81 $\times 1.00$ 7780.0 $\times 1.00$ 1.182
25 steps 24.41 $\times 2.00$ 3890.0 $\times 2.00$ 1.128
ToCa ($N=4$) 31.22 $\times 1.56$ 2540.1 $\times 3.06$ 1.111
PAB ($N=8$) 36.33 $\times 1.34$ 5962.0 $\times 1.30$ 0.8876
DuCa ($N=3,R=90\%$) 19.88 $\times 2.46$ 2873.4 $\times 2.71$ 1.152
DuCa ($N=4,R=90\%$) 15.36 $\times 3.18$ 2197.3 $\times 3.54$ 1.090
Class-conditional image generation 50,000 $256\times 256$ images uniformly sampled from the 1,000 classes in ImageNet[deng2009imagenet] are generated to evaluate the FID-50k [heusel2017fid]. Besides, sFID, Precision, and Recall are introduced as supplementary metrics.
V-A3 Results on text-to-image generation
The generation results of the proposed DuCa under various configurations are compared against recent state-of-the-art acceleration methods, including ToCa [zou2024accelerating], FORA [selvaraju2024fora], and a series of methods that directly reduce sampling steps. These comparisons are conducted on FLUX.1-dev [flux2024] and HiDream at a resolution of $1000\times 1000$, as detailed in Table II and III.
[…]
TABLE IV: Quantitative comparison on class-to-image generation on ImageNet with DiT-XL/2.
@@ V-A3 Results on text-to-image generation
[…]
DuCa(b) 0.198 $8×\times 4.59 2$.84 0.80 0.58
V-B Quantitative and Qualitative Results
Figure 4: Visualization results for different acceleration methods on FLUX.1-dev model.
[…]
V-B1 Results on text-to-video generation
The generation results of the state-of-the-art acceleration methods $\Delta$-DiT, T-GATE[tgatev1], PAB[zhao2024PAB], 15 rflow sampling steps, ToCa[zou2024accelerating], FORA [selvaraju2024fora] and proposed DuCa on the text-to-video model OpenSora[opensora] are shown in Table V.
As shown in Table V, DuCa achieves a further acceleration from 2.36 $\times$ (as with ToCa) to 2.50 $\times$ with virtually no loss in generation quality. Additionally, despite having the lowest computational cost among all the compared methods, DuCa still achieves the highest VBench score of 78.83, demonstrating that the proposed strategy in DuCa, which alternates between aggressive and conservative caching steps, effectively eliminates the redundant computations in the fully conservative caching approach.
@@ V-B1 Results on text-to-video generation
TABLE V: Quantitative comparison in text-to-video generation on VBench. *Results are from PAB [zhao2024PAB]. PAB1-3 indicate PAB with different hyper-parameters.
[…]
The advantages of DuCa, stemming from its diverse selection strategies, are also evident in its comparison with ToCa: although DuCa eliminates unnecessary redundant computations, resulting in lower actual computational load than ToCa, its VBench score decreases by only 0.58 compared to the original model, while ToCa’s score drops by 0.82, indicating that DuCa reduces quality loss by $29.3\%$.
Furthermore, thanks to DuCa’s compatibility with FlashAttention [Dao2022FlashAttentionFA], the generation of a single 480p, 2s video takes only 19.16s with DuCa, compared to 29.03s with ToCa, resulting in a speedup of nearly 10s.
@@
In Figure 6, we further present the performance of DuCa and the aforementioned methods across various dimensions of the VBench metric. DuCa shows results consistent with the non-accelerated original OpenSora model across multiple metrics. As shown in Figure 5, it is evident that DuCa is closer to the original model compared to other methods.
V-B2 Results on class-conditional image generation
@@
In Table IV, we compare ToCa[zou2024accelerating], FORA [selvaraju2024fora], the direct reduction of the corresponding DDIM steps and the proposed DuCa. The experimental results indicate that with comparable FID scores, representing generation quality, DuCa further enhances acceleration. For instance, DuCa(b) with Self-Attention score achieves a similar FID-50k as ToCa(b), but the acceleration ratio increases from 2.32 to 2.48. Additionally, under a higher acceleration scenario with an acceleration ratio around 2.7, DuCa achieves an FID-50k of 3.00, further reducing the FID-50k by 0.31 compared to ToCa’s 3.31, demonstrating that DuCa effectively reallocates computation.
Figure 6: VBench metrics and acceleration ratio of proposed DuCa and other methods.
Figure 7: Ablation study on cycle length NN and cache ratio $R$. Setting $N=5$ and $R=0.9$ yields the best trade-off between computation and quality.
V-C Ablation Studies
TABLE VI: Different caching methods on FLUX.1-dev.
Methods Image Reward FLOPs(T) $\downarrow$
Conservative-only 0.9789 1263.22
Aggressive-only 0.8539 1042.12
DuCa 0.9895 1190.24
Diversity in Token Selection
As demonstrated in Table I, the results prove that the adopted random selection method, due to its consideration of diversity in token selection, outperforms a series of traditional methods based on specific metrics while maintaining compatibility with the Efficient Attention framework. Additionally, through experiments designed around similarity, we have explored and strongly validated the importance of diversity in token selection.